# LLM-based In-situ Thought Exchanges for Critical Paper Reading

### v1

@@ Abstract.
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2025; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Human-centered computing Empirical studies in HCI
@@ 1. Introduction
A potential solution involves using conversational agents enabled by large language models (LLMs). Rather than relying solely on human peers, readers may seek out these agents as discussion partners. Previous studies have shown that interacting with LLM-enabled agents can enhance comprehension (Schmucker et al., 2024), stimulate critical thinking (Tanprasert et al., 2024), and offer multiple perspectives (Zhang et al., 2024). Recent studies further suggest that the number of agents involved differently influences people’s cognitive processing. For instance, single-agent settings are often perceived as more usable (Clarke et al., 2024), but they may help users stay engaged with a single line of thought. Alternatively, multi-agent settings more accurately simulate the dynamics of group discussion by presenting competing viewpoints (Park et al., 2023b; Zhang et al., 2024), raising counterarguments (Liang et al., 2023), and providing diverse lines of reasoning (Liu et al., 2025a). These findings suggest that both single- and multi-agent approaches could extend the intellectual benefits of peer discussion and may usefully support critical reading of academic papers.
@@ 2.1. Multi-perspective and interdisciplinary approaches in critical thinking
Critical thinking has long been recognized as a fundamental skill in education (Abrami et al., 2008; Bailin and Siegel, 2003; Dekker, 2020) and academic paper reading (Abro, [n. d.]; Coiro, 2021), encompassing higher-order abilities such as analysis, synthesis, and evaluation in Bloom’s Taxonomy (Bloom et al., 1956; Soozandehfar and Adeli, 2016; O’Riordan et al., 2021; Lee et al., 2025b).
@@
In research contexts, integrating multiple perspectives can be achieved by drawing on knowledge from different domains. Human knowledge is traditionally classified into distinct disciplines (Abbott, 2010), but there is a growing trend toward interdisciplinary study approaches (Van Noorden, 2015). In response, new tools have emerged to support interdisciplinary research and help researchers generate perspectives that integrate knowledge across domains (Zheng et al., 2024; Liu et al., 2024a, 2025b). Zheng et al. proposed DiscipLink (Zheng et al., 2024) to help researchers explore and summarize potentially relevant literature across different disciplines. Liu et al. introduced Selenite (Liu et al., 2024a), which supports readers in retrieving and reasoning about unfamiliar domain knowledge during sensemaking, enabling synthesis across multiple papers. To foster cross-disciplinary ideation, Liu et al. developed PersonaFlow (Liu et al., 2025b), which simulates experts from multiple domains to support research idea generation.
@@ 2.2. Challenges and opportunities in AI-assisted critical paper reading
Existing studies have examined various AI-supported technologies that assist academic paper reading. Most of them emphasize efficiency, helping researchers quickly extract key information (Chi et al., 2005; Fok et al., 2023, 2024; Huth et al., 2024; Gu et al., 2024) and build connections between papers via inline citations (Zhang et al., 2008; Rachatasumrit et al., 2022; Chang et al., 2023; Park et al., 2023a). However, relying on inline citations and history-based recommendations risks narrowing readers’ exposure, limiting opportunities to encounter papers from different disciplines and perspectives (Zheng et al., 2024), and potentially reinforcing echo chamber effects (Chang et al., 2023), where people are repeatedly exposed to similar opinions. Another key concern is over-reliance on AI, in which users increasingly delegate cognitive tasks to external systems, leading to cognitive offloading (Carr, 2020; Sparrow et al., 2011) and reduced active engagement in thinking. Over-reliance can negatively impact users’ independent thinking and learning processes (Gerlich, 2025; Lee et al., 2025b). For example, Gerlich et al. (Gerlich, 2025) demonstrated that reliance on AI tools undermines critical thinking performance: younger participants in their study exhibited higher dependence on AI and correspondingly lower critical thinking scores. Yatani et al. (Yatani Koji and Chi-Lan, 2024) argue that human–AI interactions should be designed to stimulate human higher-order thinking rather than encourage users to outsource cognition to AI systems. These concerns highlight a challenge in research training: junior researchers, who typically have lower domain expertise, must carefully balance leveraging AI tools with cultivating independent critical judgment.
@@
To promote critical thinking abilities, one of the most widely adopted critical reading strategies is the QRAC (Question, Read, Answer, Check) strategy (Berkeley and Riccomini, 2013; Keshav, 2007; Peng et al., 2022). It encourages readers to ask questions of the text and seek answers that deepen understanding (Wallace, 2021). Peng et al. (Peng et al., 2022) synthesized related self-questioning strategies and proposed a step-by-step guideline that supplies distinct question sets for each section of the paper reading process. They found that the proposed system enhanced junior researchers’ engagement with paper reading. Yuan et al. (Yuan et al., 2023) further improved this system by integrating questions dynamically generated by AI based on the contents of each section. Their results suggest that the revised system helped students learn to formulate critical questions. However, both systems focus narrowly on the content of a single paper, potentially constraining readers to a limited viewpoint.
Figure 1. Formative study has two sessions: (1) left is initial ideation session, participants used Google Doc to read the paper with the experimenter to observe how participants interacted with both the PDF viewer and exchanged thoughts with reading peers; (2) right is iteration session, based on the findings from the first session, we further implemented an initial prototype to gather more feedback.:
@@
Carefully-designed AI agents, however, could introduce diverse viewpoints and foster users’ critical-thinking skills. For example, SeeWidely (Zhang et al., 2024) presents a multi-agent news reading interface where users interact with diverse social agents generated from news content and their respective personas, enabling multi-perspective engagement and mitigating filter bubbles. ArgueTutor (Wambsganss et al., 2021) enables students to practice argumentation skills through adaptive, dialogue-based tutoring. Meanwhile, Tanprasert et al. (Tanprasert et al., 2024) showed that debating with chatbots adopting different personas can affect users’ critical-thinking performance in online-video-platform contexts. While prior work underscores AI’s double-edged role in supporting critical reading, it remains unclear how AI agents can actively scaffold users’ critical thinking throughout paper reading.
@@ 3.2. Experimental Design and Procedure
The formative study consisted of two sessions: the initial ideation session (Section 3.2.1) and the iteration session (Section 3.2.2). Each session consisted of three parts: a 10-minute introduction, a 30-minute paper reading task, and a 20-minute interview. The study was conducted online with Zoom.
For both sessions, we recorded and transcribed participants’ utterances for qualitative analysis.
@@ 3.2.1. Initial ideation session
Figure 2. The user interface consists of a custom PDF viewer, a left Comment Pane, and a right Section Pane. Users can highlight sentences, leave comments, generate critical thinking questions, and engage in multi-turn thought exchanges with AI agents. :
Figure 3. When users select a sentence (A) that overlaps with a pre-tagged sentence, the system automatically highlights the corresponding pre-tagged sentence (B). Users can then add the sentence to the Comment Pane and maintain the highlight in the PDF viewer (C).:
@@ 4.3. System Implementation
We implemented the system as a browser-based web application that supports both PDF rendering and interactive features. For the front-end, we used React.js  [^2] https://react.dev/ and TypeScript  [^3] https://www.typescriptlang.org/ for development, with Vite  [^4] https://vite.dev/ for project building, and the React-PDF-Viewer library  [^5] https://react-pdf-viewer.dev/ to render PDFs. For the back-end, we used Python Flask to handle the server-side logic and API requests. We used Grobid (GRO, 2025) to parse academic papers’ PDFs into structured XML files containing both content and layout information. These XML files were then provided to the OpenAI Assistants API (beta)  [^6] https://platform.openai.com/docs/assistants/migration, which supports retrieval over uploaded documents. Through this setup, the system could generate questions and feedback grounded in the paper content.
@@
Paper 16 Discussion (Liu et al., 2024b) Are these implications really insightful? How might the design implications of using AI in family expressive arts therapy inform future technological developments in human-computer interaction beyond therapeutic contexts?
@@ 5.1. Experimental Design
Multiple-agent Condition (multi condition): Previous research has suggested that interdisciplinary knowledge can enhance researchers’ performance (Van Noorden, 2015) in tasks such as literature searching (Zheng et al., 2024) and idea generation (Liu et al., 2025b). To emulate thought exchanges with researchers in different fields, we employed LLM prompts to design AI agents representing different research domains. In this condition, LLM dynamically chose three research domains based on the content of a section. Once a user pinned (Figure 2-C3) an agent, it provided feedback consistently from that disciplinary perspective throughout the interaction. The prompts used in our implementation are in Appendix D.
@@ 5.3. Procedure
The interface was deployed online using a Google Cloud Platform (GCP) Compute Engine [^8] https://cloud.google.com/compute, and participants conducted the experiments on their own personal computers. To ensure that participants did not use external generative AI tools during reading, we informed them that AI usage traces in their answers would be monitored and asked them to record their screens and submit recordings after the study. Aside from the designated reading interface, participants were only permitted to access digital libraries such as Google Scholar [^9] https://scholar.google.com, as well as translation tools like Google Translate [^10] https://translate.google.com. Eight CHI Late-Breaking Work papers (Appendix C) were selected as reading materials. All papers addressed topics related to large language models (LLMs), AI, or chatbots. These papers were relatively short and accessible, making them appropriate for junior researchers from diverse disciplinary backgrounds.
@@ 5.4.4. Perceived Workload (RQ1)
Figure 5. External review scores in five dimensions of the Critical Thinking Value Rubric, and their average as total critical thinking scores, across three groups and pre-/post-training tests.:
@@ 6.1. Response Quality Using Critical Thinking Score in Pre and Post-training Test (RQ1)
Figure 6. Average number of K-means clusters for each group at each time point (Initial vs Revised). Significance levels are indicated by asterisks: * $p<.05$, ** $p<.01$, *** $p<.001$.:
We conducted a 3 (none, single, multi; between-subject) × 2 (pre vs. post; within-subject) aligned rank transform mixed-design ANOVA on the average pre-training and post-training test scores evaluated by external reviewers. The five-dimensional critical thinking scores and total critical thinking scores across pre- and post-training tests are illustrated in Figure 5. The full results are in Appendix F.
@@
6.2. The number of unique viewpoints (RQ1)
6.3. Perceived Critical Thinking (RQ1)
Figure 7. Participants’ self-critical thinking assessment scores. Significance levels are indicated by asterisks: * $p<.05$, ** $p<.01$.:
To examine the effects of different conditions on participants’ self-perceived critical thinking ability, we conducted a Kruskal–Wallis test for average score with conditions as a factor. The results indicated a significant difference between three conditions on the total critical thinking score ($\chi^{2}(2)=8.37,p<.05,\eta^{2}=0.15$). Post-hoc pairwise comparisons using Dunn’s test with Holm correction showed that the none condition ($M=3.94,SD=1.44$) scored significantly lower than both the multi condition ($M=5.16,SD=0.97,p<.05$) and the single condition ($M=5.13,SD=0.63,p<.05$).
6.4. Perceived Workload (RQ1)
@@
6.5. Interaction Behavior Analysis (RQ2)
Figure 8. Scatter plot of participants’ interaction behaviors. The vertical axis represents the average usage counts of Reactions, and the horizontal axis represents the average usage counts of Annotations. The figure illustrates that participants in the multi condition condition relied more on reactions and less on annotations 12 out of 15 participants in multi condition never used annotations; while participants in the single condition engaged with annotations more frequently. :
6.6. Open-Ended Question Analysis (RQ2)

### v2

@@ Abstract.
copyright: acmlicensedjournalyear: 2026copyright: ccconference: 31st International Conference on Intelligent User Interfaces; March 23–26, 2026; Paphos, Cyprusbooktitle: 31st International Conference on Intelligent User Interfaces (IUI ’26), March 23–26, 2026, Paphos, Cyprusdoi: 10.1145/3742413.3789069isbn: 979-8-4007-1984-4/2026/03ccs: Human-centered computing Empirical studies in HCI
@@ 1. Introduction
A potential solution involves using conversational agents enabled by large language models (LLMs). Rather than relying solely on human peers, readers may seek out these agents as discussion partners. Previous studies have shown that interacting with LLM-enabled agents can enhance comprehension (Schmucker et al., 2024), stimulate critical thinking (Tanprasert et al., 2024), and offer multiple perspectives (Zhang et al., 2024). Recent studies further suggest that the number of agents involved differently influences people’s cognitive processing. For instance, single-agent settings are often perceived as more usable (Clarke et al., 2024), but they may keep users stay engaged with a single line of thought. Alternatively, multi-agent settings more accurately simulate the dynamics of group discussion by presenting competing viewpoints (Park et al., 2023a; Zhang et al., 2024), raising counterarguments (Liang et al., 2023), and providing diverse lines of reasoning (Liu et al., 2025b). This suggests that both single- and multi-agent approaches could extend the intellectual benefits of peer discussion and may usefully support critical reading of academic papers.
@@ 2.1. Multi-perspective and interdisciplinary approaches in critical thinking
Critical thinking has long been recognized as a fundamental skill in education (Abrami et al., 2008; Bailin and Siegel, 2003; Dekker, 2020) and academic paper reading (Abro, ; Coiro, 2021), encompassing higher-order abilities such as analysis, synthesis, and evaluation in Bloom’s Taxonomy (Bloom et al., 1956; Soozandehfar and Adeli, 2016; O’Riordan et al., 2021; Lee et al., 2025b).
@@
In research contexts, integrating multiple perspectives can be achieved by drawing on knowledge from different domains. Human knowledge is traditionally classified into distinct disciplines (Abbott, 2010), but there is a growing trend toward interdisciplinary study approaches (Van Noorden, 2015). In response, new tools have emerged to support interdisciplinary research and help researchers generate perspectives that integrate knowledge across domains (Zheng et al., 2024; Liu et al., 2024b, 2025c). Zheng et al. proposed DiscipLink (Zheng et al., 2024) to help researchers explore and summarize potentially relevant literature across different disciplines. Liu et al. introduced Selenite (Liu et al., 2024b), which supports readers in retrieving and reasoning about unfamiliar domain knowledge during sensemaking, enabling synthesis across multiple papers. To foster cross-disciplinary ideation, Liu et al. developed PersonaFlow (Liu et al., 2025c), which simulates experts from multiple domains to support research idea generation.
@@ 2.2. Challenges and opportunities in AI-assisted critical paper reading
Existing studies have examined various AI-supported technologies that assist academic paper reading. Most of them emphasize efficiency, helping researchers quickly extract key information (Chi et al., 2005; Fok et al., 2023, 2024; Huth et al., 2024; Gu et al., 2024) and build connections between papers via inline citations (Zhang et al., 2008; Rachatasumrit et al., 2022; Chang et al., 2023; Park et al., 2023b). However, relying on inline citations and history-based recommendations risks narrowing readers’ exposure, limiting opportunities to encounter papers from different disciplines and perspectives (Zheng et al., 2024), and potentially reinforcing echo chamber effects (Chang et al., 2023), where people are repeatedly exposed to similar opinions. Another key concern is over-reliance on AI, in which users increasingly delegate cognitive tasks to external systems, leading to cognitive offloading (Carr, 2020; Sparrow et al., 2011) and reduced active engagement in thinking. Over-reliance can negatively impact users’ independent thinking and learning processes (Gerlich, 2025; Lee et al., 2025b). For example, Gerlich et al. (Gerlich, 2025) demonstrated that reliance on AI tools undermines critical thinking performance: younger participants in their study exhibited higher dependence on AI and correspondingly lower critical thinking scores. Yatani et al. (Koji et al., 2024) argue that human–AI interactions should be designed to stimulate human higher-order thinking rather than encourage users to outsource cognition to AI systems. These concerns highlight a challenge in research training: junior researchers, who typically have lower domain expertise, must carefully balance leveraging AI tools with cultivating independent critical judgment.
@@
To promote critical thinking abilities, one of the most widely adopted critical reading strategies is the QRAC (Question, Read, Answer, Check) strategy (Berkeley and Riccomini, 2013; Keshav, 2007; Peng et al., 2022). It encourages readers to ask questions of the text and seek answers that deepen understanding (Wallace, 2021). Peng et al. (Peng et al., 2022) synthesized related self-questioning strategies and proposed a step-by-step guideline that supplies distinct question sets for each section of the paper reading process. They found that the proposed system enhanced junior researchers’ engagement with paper reading. Yuan et al. further improved this system by designing CriTrainer (Yuan et al., 2023), which integrates questions dynamically generated by AI based on the contents of each section. Their results suggest that the revised system helped students learn to formulate critical questions. However, existing critical reading systems primarily engage readers through a limited set of viewpoints, which may constrain readers’ exposure to alternative interpretations. Thus, to enable readers to be exposed to diverse viewpoints while reading a paper, we propose promoting thought exchanges between readers and AI agents.
@@
Indeed, carefully-designed AI agents could introduce diverse viewpoints and foster users’ critical-thinking skills. For example, SeeWidely (Zhang et al., 2024) presents a multi-agent news reading interface where users interact with diverse social agents generated from news content and their respective personas, enabling multi-perspective engagement and mitigating filter bubbles. ArgueTutor (Wambsganss et al., 2021) enables students to practice argumentation skills through adaptive, dialogue-based tutoring. Meanwhile, Tanprasert et al. (Tanprasert et al., 2024) showed that debating with chatbots adopting different personas can affect users’ critical-thinking performance in online-video-platform contexts. While prior work underscores AI’s double-edged role in supporting critical reading, it remains unclear how AI agents can actively scaffold users’ critical thinking throughout paper reading. Different from efficiency driven LLM-augmented reading tools, our work positions AI agents as discussion partners that support junior researchers’ critical reading through structured thought exchanges, emphasizing reflective engagement over direct reliance on AI-generated outputs.
@@ 3.2. Experimental Design and Procedure
Figure 1. Formative study has two sessions: (1) left is initial ideation session, participants used Google Doc to read the paper with the experimenter to observe how participants interacted with both the PDF viewer and exchanged thoughts with reading peers; (2) right is iteration session, based on the findings from the first session, we further implemented an initial prototype to gather more feedback.
An overview of the formative study setup with two sessions shown side by side. The left panel illustrates the initial ideation session, where participants read an academic paper using a Google Docs-based PDF viewer while discussing and exchanging thoughts with the experimenter. The right panel shows the iteration session, in which an early prototype was implemented based on insights from the first session and was used by participants to provide further feedback on the reading and discussion experience.
The formative study consisted of two sessions: the initial ideation session (Section 3.2.1) and the iteration session (Section 3.2.2). Each session consisted of three parts: a 10-minute introduction, a 30-minute paper reading task, and a 20-minute interview. The study was conducted online with Zoom. For both sessions, we recorded and transcribed participants’ utterances for qualitative analysis.
@@ 4. Interface Design
Figure 2. The user interface consists of a custom PDF viewer, a left Comment Pane, and a right Section Pane. Users can highlight sentences, leave comments, generate critical thinking questions, and engage in multi-turn thought exchanges with AI agents.
An overview of the system interface. The center area shows a custom PDF viewer displaying an academic paper. On the left, a Comment Pane lists user highlights and comments associated with specific sentences in the paper. On the right, a Section Pane supports generating critical thinking questions and engaging in multi-turn thought exchanges with AI agents. Users can highlight text in the PDF, leave comments, and interact with the agents in this interface.
@@ 4.1. Comment Pane
Figure 3. When users select a sentence (A) that overlaps with a pre-tagged sentence, the system automatically highlights the corresponding pre-tagged sentence (B). Users can then add the sentence to the Comment Pane and maintain the highlight in the PDF viewer (C).
An example of the sentence selection and highlighting workflow. Label A indicates a user-selected sentence in the PDF viewer. When the selected sentence overlaps with a pre-tagged sentence, label B shows that the corresponding pre-tagged sentence is automatically highlighted by the system. Label C illustrates how the selected sentence is added to the Comment Pane while the highlight is preserved in the PDF viewer.
@@ 4.2. Section Pane
4.3. Agent Design
We designed the agent personas to represent distinct disciplinary perspectives relevant to the context of the current paper, scaffolding readers’ critical reading and thought exchanges with domain-specific knowledge. To balance users’ cognitive load between reading and interacting with the agent, we designed the agents to provide responses only after users triggered events, and the agent would not engage in multi-turn conversations with the users. The prompts are in Appendix D.
@@ 4.4. System Implementation
Paper 16 Discussion (Liu et al., 2024a) Are these implications really insightful? How might the design implications of using AI in family expressive arts therapy inform future technological developments in human-computer interaction beyond therapeutic contexts?
@@
We implemented the system as a browser-based web application that supports both PDF rendering and interactive features. For the front-end, we used React.js  [^2] https://react.dev/ and TypeScript  [^3] https://www.typescriptlang.org/ for development, with Vite  [^4] https://vite.dev/ for project building, and the React-PDF-Viewer library  [^5] https://react-pdf-viewer.dev/ to render PDFs. For the back-end, we used Python Flask to handle the server-side logic and API requests. We used Grobid (24) to parse academic papers’ PDFs into structured XML files containing both content and layout information. These XML files were then provided to the OpenAI Assistants API (beta)  [^6] https://platform.openai.com/docs/assistants/migration, which supports retrieval over uploaded documents. Through this setup, the system could generate questions and feedback grounded in the paper content.
@@ 5. User Study
Figure 4. User study procedure
The User study procedure consists of four parts: (1) participants need to answer pre-study questionnaire to register the user study; (2) then participants need to complete pre-training test to read 1 paper without interface support; (3) after that, participants need to complete two weeks training periods to read another 6 papers with the interface support; (4)finally, participants need to complete post-training test to read the last paper without the interface support.
@@ 5.1. Experimental Design
Multiple-agent Condition (multi condition): Previous research has suggested that interdisciplinary knowledge can enhance researchers’ performance (Van Noorden, 2015) in tasks such as literature searching (Zheng et al., 2024) and idea generation (Liu et al., 2025c). To emulate thought exchanges with researchers in different fields, we employed LLM prompts to design AI agents representing different research domains. In this condition, LLM dynamically chose three research domains based on the content of a section. Once a user pinned (Figure 2-C3) an agent, it provided feedback consistently from that disciplinary perspective throughout the interaction. The prompts used in our implementation are in Appendix D.
@@ 5.3. Procedure
The interface was deployed online using a Google Cloud Platform (GCP) Compute Engine [^8] https://cloud.google.com/compute, and participants conducted the experiments on their own personal computers. To ensure that participants did not use external generative AI tools during reading, we informed them that AI usage traces in their answers would be monitored and asked them to record their screens and submit recordings after the study. Aside from the designated reading interface, participants were only permitted to access digital libraries such as Google Scholar [^9] https://scholar.google.com, as well as translation tools like Google Translate [^10] https://translate.google.com. Eight CHI Late-Breaking Work papers (Appendix C) were randomly selected as reading materials. All papers addressed topics related to large language models (LLMs), AI, or chatbots. These papers were relatively short and accessible, making them appropriate for junior researchers from diverse disciplinary backgrounds. In consideration of ethical concerns surrounding the upload of potentially copyrighted PDFs to cloud-based LLMs, we limited our user study materials to publicly available arXiv  [^11] https://arxiv.org/ versions.
@@ 5.4.4. Perceived Workload (RQ1)
Figure 5. External review scores in five dimensions of the Critical Thinking Value Rubric, and their average as total critical thinking scores, across three groups and pre-/post-training tests.
A quantitative comparison of external review scores using the Critical Thinking Value Rubric. Scores for five dimensions and an overall average are shown across three participant groups, with measurements taken before and after the training. The results indicate two-week training under the multi and single condition effectively enhanced critical thinking ability, together with no significant change in the none condition.
@@ 6. Results
Figure 6. Average number of K-means clusters for each group at each time point (Initial vs Revised). Significance levels are indicated by asterisks: * $p<.05$, ** $p<.01$, *** $p<.001$.
A quantitative comparison of the average number of clusters across participant groups. Values are shown for two time points, Initial and Revised, with statistical significance marked using asterisks. The results indicate that participants in single and multi condition generated more unique viewpoints than those in none condition, while there was no significant difference between single and multi condition.
Figure 7. Participants’ self-critical thinking assessment scores. Significance levels are indicated by asterisks: * $p<.05$, ** $p<.01$.
A quantitative comparison of participants’ self-assessed critical thinking scores. Results are presented across different conditions, with statistical significance marked by asterisks. Overall, Participants in both the single and multi condition reported higher self-perceived critical thinking ability than those in the none condition, with significant improvements observed across total scores as well as the Interpretation, Analysis, and Inference dimensions
@@ 6.1. Response Quality Using Critical Thinking Score in Pre and Post-training Test (RQ1)
We conducted a 3 (none, single, multi; between-subject) × 2 (pre vs. post; within-subject) aligned rank transform mixed-design ANOVA on the average pre-training and post-training test scores evaluated by external reviewers. The five-dimensional critical thinking scores and total critical thinking scores across pre- and post-training tests are illustrated in Figure 5. The full results are in Appendix F. Prior to the statistical analyses, we assessed inter-rater reliability among the four external reviewers using intraclass correlation coefficient (ICC), specifically ICC(3,k) (Shrout and Fleiss, 1979) for five dimensions. The resulting ICC values indicated moderate to good agreement across the five dimensions, ranging from 0.74 to 0.87 for the pre-training test and from 0.69 to 0.85 for the post-training test.
@@
6.2. The number of unique viewpoints (RQ1)
Figure 8. Scatter plot of participants’ interaction behaviors. The vertical axis represents the average usage counts of Reactions, and the horizontal axis represents the average usage counts of Annotations. The figure illustrates that participants in the multi condition condition relied more on reactions and less on annotations 12 out of 15 participants in multi condition never used annotations; while participants in the single condition engaged with annotations more frequently.
A scatter plot visualizing participants’ interaction behaviors. Each point represents one participant. The horizontal axis shows the average number of annotation uses, and the vertical axis shows the average number of reaction uses. Participants from different experimental conditions are distinguished using different markers or colors.
6.3. Perceived Critical Thinking (RQ1)
To examine the effects of different conditions on participants’ self-perceived critical thinking ability, we conducted a Kruskal–Wallis test for average score with conditions as a factor. The results (see Figure 7) indicated a significant difference between three conditions on the total critical thinking score ($\chi^{2}(2)=8.37,p<.05,\eta^{2}=0.15$). Post-hoc pairwise comparisons using Dunn’s test with Holm correction showed that the none condition ($M=3.94,SD=1.44$) scored significantly lower than both the multi condition ($M=5.16,SD=0.97,p<.05$) and the single condition ($M=5.13,SD=0.63,p<.05$).